Syllabus UTK – M231 – Differential Equations Fall 2005, Jochen Denzler, MWF 11:15-12:05, Ayres 214

Textbook: Nagle, Saff, Snider: Fundamentals of Differential Equations (6th ed)

We will essentially cover the following sections of the book (but minor changes are possible): 1.1–4 2.1–3,6 3.1–4 (this will go into midterm exam I) Ch. 4, with a bit of Ch. 6 spliced in (this will go into midterm exam II) Ch. 7, (this will go into midterm exam III)

All material goes in the final exam, which is scheduled by campus wide policy, as a function of the regular meeting time of the class, and cannot be rescheduled by my discretion. The time of the final exam is Fri, Dec 09, 10:15-12:15.

The midterm exams will be scheduled by me, with at least 1 week’s notice during class time. You can roughly anticipate their time from the above contents list.

Attendance: You are well advised to attend the class regularly: Textbooks at this level are not a good tool for self-study since they tend to atomize material and make you lose the big picture. The lecture will focus on compensating for this effect, and the notes I am handing out serve the same purpose. This said, I do not take formal attendance, but trust your wisdom in this matter. Be advised of the homework return policy below, which bears on the attendance issue by implication.

Homework: Be prepared to invest sufficient time in homework. Most of the hwk will come from the book, but I may add some modifications or problems not in the book. In case I schedule an occasional quiz (determined on the run), it will count towards the homework score.

The primary purpose is that you do the hwk, and that you do it well. Strive to meet the deadline. However, quality takes precedence: If I have reason to expect that extending the deadline will help you to hand in better quality (eg, if questions have arisen and you want to incorporate feedback I give you about your questions), I am prepared to give an extension.

Graded homework will be returned in class (or, if you visit the office hour, you may also request it there). Any particular homework which I have not been able to return to its author for three subsequent class sessions will be \underline{discounted} 50%, unless there is a good reason (like e.g., extended sickness, university travel). !!! The purpose of this policy is that an individual who ostensibly doesn’t care about the feedback on his/her homework wastes good resources and is likely to make the same mistakes over again. Learning from the feedback is part of the homework.

Your course grade will be determined as

– either homework and 3 midterm exams: each 1/6; plus 1/3 final exam

– or homework 1/6; plus 3 midterm exams each 1/9; plus 1/2 final exam, whichever gives the better point score for you.

You are guaranteed an A with 90%, a B with 80%, a C with 70%, a D with 55%. However exam difficulties may vary, and I will set the precise cutoff points at the end; they may be a bit more lenient than the minimum guarantee.

Be sure not to procrastinate difficulties and ask questions of understanding as soon as they arise. This course should have the side effect of making calculus more meaningful. It should therefore not be reduced to a formula crunching course.

Office: I have tentatively scheduled regular office hours MWF 10:10-11:00, but we can later switch to other times if these are inconvenient for you. You are also welcome to request an appointment at other times.

I accomodate drop-ins, whenever feasible, even though I cannot always guarantee immediate availability. My office is Ayres 317 E, phone 4-5325. If you have difficulty reaching me otherwise, you may try my cell phone 604-7173. I want to be available for questions, whenever possible, even though I may not always be in the office. Email is denzler@math.utk.edu, but I may not read it for half a day or an entire weekend.

— There is one exception: I am likely not in and/or not available on Tuesdays, which I reserve as research priority day.

A course web page is set up from my home page:

http://www.math.utk.edu/~denzler/M231-Fa2005/.

Bookmark it. I will use the BlackBoard system to get updated class lists, and possibly for email broadcasts, but material of general interest will be posted on the public webpage off my homepage.